OpenStudy (anonymous):

Let f(x)= -2x+4 and g(x)= -6x-7 find f(x)-g(x). Please help

OpenStudy (triciaal):

f(x) - g(x) = -2x + 4 - (-6x -7)= -2x + 4 + 6x + 7 =

OpenStudy (anonymous):

i got 4x+11?

OpenStudy (triciaal):

yes 4x + 11

OpenStudy (anonymous):

can u help me with some more?????

OpenStudy (triciaal):

post

OpenStudy (anonymous):

f(x)=3x-7 and g(x)=-2x-6 find (f o g)(4)

OpenStudy (triciaal):

follow the same process notice I use the parenthesis

OpenStudy (anonymous):

oh ok

OpenStudy (anonymous):

What is the inverse of the given relation??? y=3x+12

OpenStudy (triciaal):

fog means use the value of g in f(x) find the value when x = 4 g(x)=-2x-6 and x = 4 g(4) = -2(4) -6 = -8-6 = -14 f(-14) = 3(-14)-7=-42 -7 = -49 3(-2x-6) -7 = -6x -18 -7 = -6x -25 x = 4 so -6*4 - 25 = -49 notice it does not matter if you work with x first or if you substitute first

OpenStudy (triciaal):

the inverse of the relation y = whatever with x is to make x the subject so x = whatever with y in it.

OpenStudy (triciaal):

y=3x+12 find x =

OpenStudy (anonymous):

oh ok!